An insurance fund invests Rs. 500,000 at the beginning of each year for five years at an interest rate of 8.5% p.a. Calculate the total amount in the fund at the end of year 5.


1
Expert's answer
2021-11-15T16:48:08-0500

Amount invested each year = P = $500000

Number of years = n = 5

Interest Rate = r = 8.5%

Hence, Future Value "= P[\frac{((1+r)^n -1)}{r}](1+r)"

= "500000[\frac{((1+0.085)^5 -1)}{0.085}](1+0.085)" = $3,214,514.75